# On System Operators with Variation Bounding Properties

### v2

@@ 3.2 The $\text{VB}_{k-1}$ case
Proposition 7 is a result of Lemma 3 in conjunction with the following characterization of strictly totally positive matrices (see, e.g., [6, Theorem 2.1]).
@@
Let $\in\mathds{R}^{n\times m}. XX$ is strictly totally positive if and only if all row and column initial minors of $X$ are positive.
For completion, a new proof of Proposition 10, similar to [28, Theorem 2.3], is stated in the appendix. Based on that, we show the following new non-strict relaxation similar to Proposition 6.
@@
Let $X\in\mathds{R}^{n\times m}$ be such that row and column initial $k$-order minors of $X$ are positive for $k=1,2,...,\min\{n-1,m-1\}$ and nonnegative for $k=\min\{n,m\}$. Then, $X$ is totally positive and all minors of order less than $k$ are positive.
@@
with $P_{\beta}:={I}_{(1:n),\beta}$. Therefore, $\mathcal{O}$} is $\text{VB}_{k-1}$ if $(\tilde{A}_{r},\tilde{b}_{k,r,\beta},\tilde{c}_{r})$ is strictly externally positive/negative for all $(r,\beta)$ with $r\in(1:k)$ and $\beta$ as in Corollary 1, except for the following cases $\beta=(\bar{t}:\bar{t}+\bar{r}-1)$, $\bar{r}=k$, $\bar{t}\in(k+1:m-k+1)$, where non-strict externally positive/negative is sufficient if $\tilde{c}_{r}\tilde{A}_{[r]}^{l-1}\tilde{b}_{k,r,\beta}\neq 0$ for all $l\in(1:k)$.
@@ 3.3 The $\text{OVD}_{k}/\penalty 50\text{VD}_{k}$ case
Finally, for $\text{VD}_{k}$ there exists an equivalence to ${SR}_{k}$ with a full rank requirement [19, Theorem 5.1.4]. Based on Theorem 4, we are able to relax this requirement, too.
@@ 4 Discussion
    The number of under- and overshoots in a step response, which may be used as measure for the trade-off between fast rise vs. fast settling time, equals the number of sign changes of the impulse response. Yet, mathematically, little is known about the tractability of this measure. While there exist several lower bounds [3, 31, 5], aside from external positivity certificates [altafini2016minimal, 15, vandeneijnden2025externally, 32, 8], only few methods provide generic upper bounds [5, 14]. By noticing that the impulse response of (1) can be expressed as $g={\mathcal{O}}(A,c)b$, our results can be used to bound the number of sign changes in $g$ via
@@
    where $\|u\|_{\ell_{\infty}}:=\max_{t\in(0:N-1)}|u(t)|$, are known to lead (e.g., by the alignment property [20, Chapter 5]) to solutions of the form
@@
    for some $\mu\in\mathds{R}^{n}$, $u_{\max}>0$, and otherwise $u^{\ast}$ may take any value in $u_{max},u_{\max}]$. In other words, the number of sign changes (bangs) of $u^{\ast}$ are bounded by the strict variation of the impulse response $g_{\mu}(t):=\mu^{\mathsf{T}}A^{-t}A^{-1}b$. Using Theorem 2, we can conclude similar to (4) that if ${\mathcal{C}}(A^{-1},A^{-1}b)$ is $\text{SVB}_{k}$ and $\text{S}^{-}(\mu)\leq k$, then $u^{\ast}$ has at most $k$ bangs. This is an extension and discrete-time analog of the well-known continuous-time characterization, which asserts that if $(A,b)$ is controllable and $\sigma(A)\subset\mathds{R}$, then $u^{\ast}$ has at most n1n-1 bangs [11].
Further, note that Theorems 4 and 5 as well as Propositions 9 and 12 haven been instrumental for new tractable approaches in analyzing basis pursuit in structured problems, incl. fuel optimal control [22], as well as the study of self-oscillations in relay feedback systems [34, 13].
@@ 5 Illustrative Examples
In this section, we want to illustrate our results based on three system examples (1) for which we want to bound the variation of the their impulse responses via the relationship $g={\mathcal{O}}b$. The three systems have the following distinct properties:
@@
In all examples, we employ [15] to check external positivity/negativity of the impulse response as required by Theorem 2 and Corollary 4.
@@ 5.3 Example 3: $\text{SVB}_{1}$ $\mathcal{O}$} without $k$-positive AA representation
can be shown to be negative. The operator resulting from reversing the column order in $\bar{\mathcal{H}}_{g}$ then has positive $2$-minors and as such fulfills the requirements of Proposition 6. As ${SC}_{2}$ is invariant to column inversion, this shows that $\bar{\mathcal{H}}_{g}$ is ${SR}_{2}$ and by Corollary 5, we conclude that $\mathcal{O}$} is $\text{VD}_{1}$. Finally, as the dominant eigenvalues in $\bar{A}$ include complex ones with arguments that are non-rational multiples of $\pi$, it can can be shown by Lemma 1 and [1, Theorem 5] that there does not exist any realization where $A$ is ${SC}_{2}$.
@@ 6 Conclusion
We have derived a tractable approach to certify that the observability operator of a discrete-time LTI system maps a vector with at most $k$ sign changes to a sequence with at most $k$ sign changes. Our derivations are based on new complements to the existing matrix literature of $k$-variation bounding matrices and their application to the observability operator. These complements include the relaxation of rank and dimension assumptions for the characterization of $k$-variation bounding matrices as well as their simplified certification. Interestingly, application of those results to the observability operator transform our problem into a problem of externally positivity verification. While this generalizes recent certificates for the more restrictive notion of order-preserving $k$-variation diminishment [14], it became also evident that variation bounding and variation diminishing properties in system operators share the same eigenvalue requirements, i.e., the $k+1$ largest eigenvalues of $A$ have to be real and positive.
An advantage of our certificate over [14] is the removal of the $k$-positivity requirement on $A$. In particular, we have constructed an example of a variation-diminishing observability operator, where independent of the system realization such an $A$ does not exist. Our results are applied to the problem of bounding the number of sign changes in the impulse response.
@@ Appendix A Proof to Proposition 3
We only prove Items iii & v, as the others follow from [19, Propositions 5.1.1 & 5.1.2] and [14, Propositions 2.7].
@@ Appendix C Proof to Proposition 9
In order to conclude that $X$ is ${SSC}_{k}$, we need to show strict sign agreement between all the $X_{(1:n),\alpha}$. By (6), it suffices to show this only for the first $k$ columns, i.e., that $X_{(1:k),\beta}$ is ${SSC}_{k}$ for all $\beta\in\mathcal{I}_{m,k+1}$. To this end, let $\bar{u}\in\mathds{R}^{k+1}$ be such that ${X}_{(1:k),\beta}\bar{u}=0$ and $X_{(1:k),\beta_{(1:k)}}\bar{u}_{(1:k)}=X_{(1:k),\beta_{\{k+1\}}}$. Then, by Cramer’s rule [18, Sec. 0.8.3],
@@
Assuming the contrary, there exits at least one $j^{\ast}$ such that $\bar{u}_{\beta_{j^{\ast}}}\bar{u}_{\beta{j^{\ast}+1}}>0$. Then, $\text{S}^{-}(\bar{u})\leq k-1$ and $S^{+}(X_{(1:n),\beta}\bar{u})\geq k$, as the first $k$ elements in $X_{(1:n),\beta}\bar{u}$ are zero. This is a contradiction, since $X_{(1:n),\beta}$ is $\text{SVB}_{k-1}$ by assumption.
@@ Appendix E Proof to Corollary 2
We begin by noting that by Item ii in Lemma 1, the dominant eigenvalue of $A_{[r]}$ is given by $\prod_{i=1}^{r}\lambda_{i}(A)$. Further, since by assumption $\mathcal{O}$} is ${SSC}_{k}$, it follows from Theorem 2 that for each $\beta\in\mathcal{I}_{n,k}$, the impulse response of the systems defined in Theorem 2 have to be strictly externally positive/negative. Since this requires that each system has a dominant positive pole [24, 8, 1], it follows that $\prod_{i=1}^{r}\lambda_{i}(A)\neq 0$ and our claim will be proven by Theorem 2 if we can show that $\prod_{i=1}^{r}\lambda_{i}(A)$ is a controllable and observable mode of $(\tilde{A}_{r},\tilde{b}_{k,r,\beta},\tilde{c}_{r})$ for each $r\in(1:k)$ and at least one corresponding choice of $\beta\in\mathcal{I}_{n,k}$.
@@
and [18, Sec. 0.9.11] that all elements in $\bar{c}_{r}$ are non-zero. Further, the first element in $\bar{b}_{k,r,\beta}$ is zero for all $\beta\in\mathcal{I}_{n,k}$ if and only if
@@
because $|v_{1}|=|\det(\bar{\mathcal{O}}^{n}_{(1:k-r,r+1:k)})|\neq 0$ by [18, Sec. 0.9.11]. Hence, there exists at least one $\beta$ such that the first element $\bar{b}_{r}$ is non-zero. This concludes the proof for diagonalizable $A$. Finally, based on the previous case, we will show, now, that in general the eigenvalues $\lambda_{1}(A),\dots,\lambda_{k}(A)$ are simple. To this end, assume the opposite, i.e., at least one of the $k$ dominant eigenvalues has algebraic multiplicity greater than 1. Then, by the observability of $(A,c)$, AA must be non-diagonalizable (e.g., from the Popov-Belevitch-Hautus test [11]) such that there exists a $T\in\mathds{C}^{n\times n}$ with
@@ Appendix G Proof to Theorem 4
    It remains to show that each ${\left({X}_{(1:n),I}\right)}_{[k]}$ consists of elements that share the same sign accross all possible choices of $I$. This will be shown by considering a sequence of submatrices with properties that interlink all ${\left({X}_{(1:n),I}\right)}_{[k]}$. To this end, let $J^{*}\in\mathcal{I}_{m,k+1}$ be a fixed index set such that $\textnormal{rank}\left({X}_{(1:n),J^{*}}\right)=k+1$ ($J^{\ast}$ always exists because $\textnormal{rank}(X)>r$). The sequence $\{{X}_{(1:n),J^{i}}\}_{i=1}^{k+1}$ of submatrices in $\mathds{R}^{n\times k+1}$ is then defined by a sequence of index sets $\{J^{i}\}_{i=1}^{k+1},\;J^{i}\in\mathcal{I}_{m,k+1}$ with the following properties:
@@
    Next $J^{2}$ has one element from $I$ replaced with an element from $J^{*}$ such that $\textnormal{rank}\left({X}_{(1:n),J^{2}}\right)=k+1$. This is repeated until the sequence is complete. In the case that elements of $I$ overlap with elements of $J^{*}$, the sequence will be created using the same steps with less elements in total.
@@ Appendix I Proof to Proposition 11
The proof is similar to the proof in [26, Theorem 2.2(i)]. The idea is to use the bijection in Lemma 3 between the original matrix $X$ and another matrix, $C$, which has to be totally positive for $X$ to be ${SC}_{m}$. Using our modification for the non-strict case in Proposition 11, this switches a strict sign requirement to a non-strict sign requirement for some of the minors in $C$ and, hence, $X$. In particular, the minors in Lemma 3 that are allowed to be non-strict are those with
@@
From the proof of Theorem 2 the equivalence between the impulse responses and the minors is already established. Corollary 3 can be used then to establish the above relaxed requirement of (strict) external positvity/negativity of the system corresponding the non-strict sign consistency assumption.

### v3

@@ 3.2 The $\text{VB}_{k-1}$ case
Proposition 7 is a result of Lemma 3 in conjunction with the following characterization of strictly totally positive matrices (see, e.g., [fallat2017total, Theorem 2.1]).
@@
$X\in\mathds{R}^{n\times m}$ is strictly totally positive if and only if all row and column initial minors of $X$ are positive.
For completion, a new proof of Proposition 10, similar to [pinkus2009totally, Theorem 2.3], is stated in the appendix. Based on that, we show the following new non-strict relaxation similar to Proposition 6.
@@
Let $X\in\mathds{R}^{n\times m}$ be such that row and column initial $k$-order minors of $X$ are positive for $k=1,2,...,\min\{n-1,m-1\}$ and nonnegative for $k=\min\{n,m\}$. Then, $X$ is totally positive and all minors of order less than $\min\{n,m\}$ are positive.
@@
with $P_{\beta}:={I}_{(1:n),\beta}$. Therefore, $\mathcal{O}$} is $\text{VB}_{k-1}$ if $(\tilde{A}_{r},\tilde{b}_{k,r,\beta},\tilde{c}_{r})$ is strictly externally positive/negative for all $(r,\beta)$ with $r\in(1:k)$ and $\beta$ as in Corollary 1, except for the following cases $\beta=(\bar{t}:\bar{t}+\bar{r}-1)$, $\bar{r}=k$, $\bar{t}\in(k+1:m-k+1)$, where non-strict external positivity/negativity is sufficient if $\tilde{c}_{r}\tilde{A}_{[r]}^{l-1}\tilde{b}_{k,r,\beta}\neq 0$ for all $l\in(1:k)$.
@@ 3.3 The $\text{OVD}_{k}/\penalty 50\text{VD}_{k}$ case
Finally, for $\text{VD}_{k}$ there exists an equivalence to ${SR}_{k}$ with a full rank requirement [karlin1968total, Theorem 5.1.4]. Based on Theorem 4, we are able to relax this requirement, too.
@@ 4 Discussion
    The number of under- and overshoots in a step response, which may be used as measure for the trade-off between fast rise vs. fast settling time, equals the number of sign changes of the impulse response. Yet, mathematically, little is known about the tractability of this measure. While there exist several lower bounds [damm2014zero, swaroop1996some, elkhoury1993discrete], aside from external positivity certificates [altafini2016minimal, grussler2019tractable, vandeneijnden2025externally, taghavian2023external, farina2011positive], only few methods provide generic upper bounds [elkhoury1993discrete, grussler2021internally]. By noticing that the impulse response of (1) can be expressed as $g={\mathcal{O}}(A,c)b$, our results can be used to bound the number of sign changes in $g$ via
@@
    where $\|u\|_{\ell_{\infty}}:=\max_{t\in(0:N-1)}|u(t)|$, are known to lead (e.g., by the alignment property [luenberger1968optimization, Chapter 5]) to solutions of the form
@@
    for some $\mu\in\mathds{R}^{n}$, $u_{\max}>0$. Otherwise, $u^{\ast}(t)$ may take any value in $u_{max},u_{\max}]$. In other words, the number of sign changes (often called bangs) of $u^{\ast}$ are bounded by the strict variation of the impulse response $g_{\mu}(t):=\mu^{\mathsf{T}}A^{-t}A^{-1}b$. Using Theorem 2, we can conclude similar to (4) that if ${\mathcal{C}}(A^{-1},A^{-1}b)$ is $\text{SVB}_{k}$ and $\text{S}^{-}(\mu)\leq k$, then $u^{\ast}$ has at most $k$ bangs. This is an extension and discrete-time analog of the well-known continuous-time characterization, which asserts that if $(A,b)$ is controllable and $\sigma(A)\subset\mathds{R}$, then $u^{\ast}$ has at most n1n-1 bangs [glad2000control].
Further, note that Theorems 4 and 5 as well as Propositions 9 and 12 have been instrumental for new tractable approaches in analyzing basis pursuit in structured problems, incl. fuel optimal control [marmary2025tractable], as well as the study of self-oscillations in relay feedback systems [tong2025selfsustained, grussler2025discrete].
@@ 5 Illustrative Examples
In this section, we want to illustrate our results based on three system examples (1) for which we will bound the variation of their impulse responses via the relationship $g={\mathcal{O}}b$. The three systems have the following distinct properties:
@@
In all examples, we employ [grussler2019tractable] to check external positivity/negativity of the impulse response as required by Theorem 2 and Corollary 4.
@@ 5.3 Example 3: $\text{SVB}_{1}$ $\mathcal{O}$} without $k$-positive AA representation
can be shown to be negative. The operator resulting from reversing the column order in $\bar{\mathcal{H}}_{g}$ then has positive $2$-minors and as such fulfills the requirements of Proposition 6. As ${SC}_{2}$ is invariant to column inversion, this shows that $\bar{\mathcal{H}}_{g}$ is ${SR}_{2}$ and by Corollary 5, we conclude that $\mathcal{O}$} is $\text{VD}_{1}$. Finally, as the dominant eigenvalues in $\bar{A}$ include complex ones with arguments that are non-rational multiples of $\pi$, it can can be shown by Lemma 1 and [benvenuti2004tutorial, Theorem 5] that there does not exist any realization where $A$ is ${SC}_{2}$.
@@ 6 Conclusion
We have derived a tractable approach to certify that the observability operator of a discrete-time LTI system maps a vector with at most $k$ sign changes to a sequence with at most $k$ sign changes. Our derivations are based on new complements to the existing matrix literature of $k$-variation bounding matrices and their application to the observability operator. These complements include the relaxation of rank and dimension assumptions for the characterization of $k$-variation bounding matrices as well as their simplified certification. Interestingly, application of those results to the observability operator transform our problem into a problem of externally positivity verification. While this generalizes recent certificates for the more restrictive notion of order-preserving $k$-variation diminishment [grussler2021internally], it became also evident that variation bounding and variation diminishing properties in system operators share the same eigenvalue requirements, i.e., the $k+1$ largest eigenvalues of $A$ have to be real and positive.
An advantage of our certificate over [grussler2021internally] is the removal of the $k$-positivity requirement on $A$. In particular, we have constructed an example of a variation-diminishing observability operator, where independent of the system realization such an $A$ does not exist. Our results are applied to the problem of bounding the number of sign changes in the impulse response.
@@ Appendix A Proof to Proposition 3
We only prove Items iii & v, as the others follow from [karlin1968total, Propositions 5.1.1 & 5.1.2] and [grussler2021internally, Propositions 2.7].
@@ Appendix C Proof to Proposition 9
In order to conclude that $X$ is ${SSC}_{k}$, we need to show strict sign agreement between all the $X_{(1:n),\alpha}$. By (6), it suffices to show this only for the first $k$ columns, i.e., that $X_{(1:k),\beta}$ is ${SSC}_{k}$ for all $\beta\in\mathcal{I}_{m,k+1}$. To this end, let $\bar{u}\in\mathds{R}^{k+1}$ be such that ${X}_{(1:k),\beta}\bar{u}=0$ and $X_{(1:k),\beta_{(1:k)}}\bar{u}_{(1:k)}=X_{(1:k),\beta_{\{k+1\}}}$. Then, by Cramer’s rule [horn2012matrix, Sec. 0.8.3],
@@
Assuming the contrary, there exists at least one $j^{\ast}$ such that $\bar{u}_{\beta_{j^{\ast}}}\bar{u}_{\beta{j^{\ast}+1}}>0$. Then, $\text{S}^{-}(\bar{u})\leq k-1$ and $S^{+}(X_{(1:n),\beta}\bar{u})\geq k$, as the first $k$ elements in $X_{(1:n),\beta}\bar{u}$ are zero. This is a contradiction, since $X_{(1:n),\beta}$ is $\text{SVB}_{k-1}$ by assumption.
@@ Appendix E Proof to Corollary 2
We begin by noting that by Item ii in Lemma 1, the dominant eigenvalue of $A_{[r]}$ is given by $\prod_{i=1}^{r}\lambda_{i}(A)$. Further, since by assumption $\mathcal{O}$} is ${SSC}_{k}$, it follows from Theorem 2 that for each $\beta\in\mathcal{I}_{n,k}$, the impulse response of the systems defined in Theorem 2 have to be strictly externally positive/negative. Since this requires that each system has a dominant positive pole [ohta1984reachability, farina2011positive, benvenuti2004tutorial], it follows that $\prod_{i=1}^{r}\lambda_{i}(A)\neq 0$ and our claim will be proven by Theorem 2 if we can show that $\prod_{i=1}^{r}\lambda_{i}(A)$ is a controllable and observable mode of $(\tilde{A}_{r},\tilde{b}_{k,r,\beta},\tilde{c}_{r})$ for each $r\in(1:k)$ and at least one corresponding choice of $\beta\in\mathcal{I}_{n,k}$.
@@
and [horn2012matrix, Sec. 0.9.11] that all elements in $\bar{c}_{r}$ are non-zero. Further, the first element in $\bar{b}_{k,r,\beta}$ is zero for all $\beta\in\mathcal{I}_{n,k}$ if and only if
@@
because $|v_{1}|=|\det(\bar{\mathcal{O}}^{n}_{(1:k-r,r+1:k)})|\neq 0$ by [horn2012matrix, Sec. 0.9.11]. Hence, there exists at least one $\beta$ such that the first element $\bar{b}_{r}$ is non-zero. This concludes the proof for diagonalizable $A$. Finally, based on the previous case, we will show, now, that in general the eigenvalues $\lambda_{1}(A),\dots,\lambda_{k}(A)$ are simple. To this end, assume the opposite, i.e., at least one of the $k$ dominant eigenvalues has algebraic multiplicity greater than 1. Then, by the observability of $(A,c)$, AA must be non-diagonalizable (e.g., from the Popov-Belevitch-Hautus test [glad2000control]) such that there exists a $T\in\mathds{C}^{n\times n}$ with
@@ Appendix G Proof to Theorem 4
    It remains to show that each ${\left({X}_{(1:n),I}\right)}_{[k]}$ consists of elements that share the same sign across all possible choices of $I$. This will be shown by considering a sequence of submatrices with properties that interlink all ${\left({X}_{(1:n),I}\right)}_{[k]}$. To this end, let $J^{*}\in\mathcal{I}_{m,k+1}$ be a fixed index set such that $\textnormal{rank}\left({X}_{(1:n),J^{*}}\right)=k+1$ ($J^{\ast}$ always exists because $\textnormal{rank}(X)>r$). The sequence $\{{X}_{(1:n),J^{i}}\}_{i=1}^{k+1}$ of submatrices in $\mathds{R}^{n\times k+1}$ is then defined by a sequence of index sets $\{J^{i}\}_{i=1}^{k+1},\;J^{i}\in\mathcal{I}_{m,k+1}$ with the following properties:
@@
    Next $J^{2}$ has one element from $I$ replaced with an element from $J^{*}$ such that $\textnormal{rank}\left({X}_{(1:n),J^{2}}\right)=k+1$. This is repeated until the sequence is complete. In the case that elements of $I$ overlap with elements of $J^{*}$, the sequence will be created using the same steps with fewer elements in total.
@@ Appendix I Proof to Proposition 11
The proof is similar to the proof in [pena_matrices_1995, Theorem 2.2(i)]. The idea is to use the bijection in Lemma 3 between the original matrix $X$ and another matrix, $C$, which has to be totally positive for $X$ to be ${SC}_{m}$. Using our modification for the non-strict case in Proposition 11, this switches a strict sign requirement to a non-strict sign requirement for some of the minors in $C$ and, hence, $X$. In particular, the minors in Lemma 3 that are allowed to be non-strict are those with
@@
From the proof of Theorem 2 the equivalence between the impulse responses and the minors is already established. Corollary 3 can be used then to establish the above relaxed requirement of (strict) external positivity/negativity of the system corresponding the non-strict sign consistency assumption.